OpenStudy (anonymous):

Find the volume of a prism with a square base that is 5 cm by 5 cm and is 10 cm tall.

jimthompson5910 (jim_thompson5910):

V = LWH V = 5*5*10 V = 25*10 V = 250 So the volume is 250 cubic cm